OpenStudy (anonymous):

find the general solution of the differential equation y' = 3x^2sqrt(1-y^2)

OpenStudy (saifoo.khan):

sorry sonofa, i forgot calculus otherwise i must have helped.

OpenStudy (anonymous):

dy/dx=3x^2sqrt(1-y^2) dy/sqrt(1-y^2)=3x^2dx integrate both sides sin^-1(y)=x^3+C y=sin(x^3+C)